1. A mixture of gases consists of 40 percent argon and 60 percent nitrogen by mass. The mixture is expanded from 250 kPa and 400 C to 150 kPa and 250°C in an adiabatic, steady-flow turbine. Determine the work of the turbine. State the assumptions and show the complete reductions of the governing equations.
1
Expert's answer
2021-10-28T02:59:49-0400

Molar mass of Argon= 39.948

0.4×39.948= 15.98g

Molar mass of Nitrogen = 14.0067

14.0067×0.6= 8.4g

250-400= -150

150-250= -100

= -150×-100×24.38= 365.7kJ